Two bicycle tires are set rolling with the same initial speed of 3.10 m/s along a long, straight road, and the distance each travels before its speed is reduced by half is measured. One tire is inflated to a pressure of 40 psi and goes a distance of 19.0 m ; the other is at 105 psi and goes a distance of 92.5 m . Assume that the net horizontal force is due to rolling friction only and take the free-fall acceleration to be g = 9.80 m/s2 .

What is the coefficient of rolling friction μr for the tire under low pressure?

μr =___________________________
Answer #1

Vi = 3.10 m/s

Vf=Vi/2

=3.10/2
Vf = 1.55 m/s
s = 19.0 m

For 40 psi pressure -
Vf^2 = Vi^2 - 2*a*s
1.55^2 = 3.10^2 - 2*a*19
a1 = 0.190 m/s^2

For 105 psi pressure -
Vf^3 = Vi^2 - 2*a*s
1.55^2 = 3.10^2- 2*a*92.5
a2 = 0.0389 m/s^2

Friction Force, Fs = μr*m*g
m*a = μr*m*g
a = μr*g
μr = a/g

Coefficient of friction for 40 psi pressure tyre -
μr = 0.190 / 9.8
μr = 0.0194

Similarly Coefficient of friction for 105 psi pressure tyre -
μr = 0.0389 / 9.8
μr = 0.00397

I hope help you !!